OpenStudy (confusedhampster):

Math has always been really hard for me. I always need help, and I have no one around the house to help me. Please help! Example A: 2n + 3 = 13 Example B: n + 3 = 13 Which example is a one-step equation? How is solving a two-step equation like solving one-step equation? How is solving them different?

OpenStudy (imstuck):

B is a one-step equation, because it only takes one step to solve it for n. Your one step would be to subtract 3 from both sides and you have the equation in terms of n.

OpenStudy (imstuck):

Solving a two-step is like solving a one step because the end result will be the same...putting the equation in terms of a certain variable. They are different because two-step equations often take some mathematical "gymnastics" to get the desired variable alone. Like taking a square root of a number or dividing it or multiplying it...the list is endless.

OpenStudy (anonymous):

n=13-3/2=5, n=13-3=10

OpenStudy (confusedhampster):

I have one more: Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right. How can I answer this if I dont even understand the lesson! :( HELP PLEASE!!!!!!!!!!!!!!!

OpenStudy (anonymous):

wait